# AI-driven Conservative-to-Primitive Conversion in Hybrid Piecewise Polytropic and Tabulated Equations of State

### v1

@@ II.3 Training approach
$\mathcal{L}(\theta)=\frac{1}{n}\sum_{i=1}^{n}(\hat{y}_{i}(\theta)-y_{i})^{2}+q\cdot\sum_{i=1}^{n}\text{ReLU}(-\mathcal{N}^{-1}(\hat{y}_{i}(\theta))),$
where $\hat{y}_{i}(\theta)$ represents the network’s estimation for feature $i$, $y_{i}$ is the corresponding target value, ReLU is the familiar rectified linear unit defined by $\text{ReLU}(x)=\max(0,x)$, and $\mathcal{N}^{-1}(\cdot)$ represents an inverse normalization procedure based on the training data statistics. The penalty factor, $q$, was optimized for each model, with $q=150$ for NNC2PS and $q=350$ for NNC2PL. These values consistently suppressed negative pressure predictions on the test set. For the tabulated EOS model (NNC2P_Tabulated), the structure of the data precluded negative predictions, so a standard mean-squared error loss function was used.
@@
II.4 Inference Speed Tests
@@ II.4.1 TorchScript Deployment
II.4.2 TensorRT Deployment
For TensorRT, we explored both FP32 (unquantized) and FP16-quantized engines, ultimately deciding not to pursue INT8 quantization due to accuracy degradation observed in initial tests. After extensive testing, we opted for dynamic engine building with batch sizes up to approximately 25,000, as this configuration provided the best balance between performance and flexibility for our hardware and model structure. It must be noted that constructing an optimal engine in TensorRT is a nuanced process, influenced by multiple factors including model architecture, hardware specifications, intended batch sizes during inference, and input data. Therefore, achieving the best results often involves iterative tuning and profiling to adapt the engine to the specific deployment environment and workload requirements. Below, we summarize the overall engine-building process we followed in detail:
@@ II.4.2 TensorRT Deployment
  • Configuration and Optimization Profiles: The BuilderConfig was set with a 2 GB workspace memory limit, balancing memory usage and speed. We set up a dynamic optimization profile specifying minimum, optimal, and maximum batch sizes, tailored to accommodate varying input sizes efficiently. This profile allowed the engine to handle batch sizes dynamically, with an optimal batch size close to our expected usage, while allowing flexibility to adapt for both smaller and larger input volumes.
@@
For the actual inference speed test procedure, we implemented two distinct workflows on a single GPU for both approaches. The TorchScript-based approach required no special configuration beyond loading the model and executing it in batches. It utilized libtorch to load the model directly, allowing straightforward GPU deployment and batch processing.
For the TensorRT-based approach, however, several additional configurations were required. The model, after conversion to an optimized engine, was loaded using TensorRT’s C++ API. This process required close attention to low-level CUDA memory management through TensorRT’s executeV2 function, setting up input and output buffers manually, and optimizing precision settings (such as enabling half-precision FP16) to enhance performance. While TensorRT’s setup was more complex, it promises to enable significant inference speed improvements by utilizing hardware-specific optimizations.
III Results
III.0.1 Accuracy
@@
Figure 4: Inference speed comparison of various C2P inversion methods. (a) Inference time as a function of dataset size for a traditional numerical solver (RePrimAnd on a single CPU) and two neural network models (NNC2PS and NNC2PL) using TensorRT (FP32 and FP16 precision) and TorchScript on a single NVIDIA A100 GPU. (b) Inference speed comparison for a 200,000 data point dataset, highlighting the performance gains achieved with TensorRT, particularly with FP16 optimization.
@@
III.0.2 Inference Speed Analysis
@@
The inference speed of various approaches was analyzed using a single NVIDIA A100 GPU for neural network models and a CPU implementation of the numerical method, as in the RePrimAnd library, to represent traditional simulation techniques. Each setup was tested across multiple dataset sizes ranging from 25,000 to 200,000 data points, with ten inference runs conducted per configuration to ensure stability and consistency in the results, as can be seen in Fig. 4(a).
RePrimAnd-based implementation of the numerical method executed all computations on a CPU and demonstrated nearly linear scaling with respect to the size of the dataset. While it performed adequately for smaller datasets, the approach became increasingly time-consuming for larger datasets. For 200,000 datapoints, for example, it was almost 8 times slower than the fastest TensorRT engine (See Fig. 4(b)).
The neural network-based methods NNC2PS and NNC2PL were deployed in three configurations: using TensorRT with both FP32 and FP16 precision modes, and utilizing TorchScript format. All configurations utilized the NVIDIA A100 GPU for acceleration. Neural networks deployed in full precision with TensorRT consistently (and expectedly) delivered faster inference times than the numerical method, achieving substantial improvements even for the largest datasets. However, the computational requirements of full precision meant that the performance gains were modest compared to the optimized FP16 configuration. Mixed-precision deployments in TensorRT offered the most significant performance gains across all dataset sizes. In this configuration, both NNC2PS and NNC2PL achieved substantial reductions in inference time compared to full precision. The FP16 optimizations allowed these models to process datasets ranging from 25,000 to 200,000 data points with remarkable efficiency, showcasing their suitability for large-scale applications. Finally, as a general-purpose deployment pipeline, TorchScript demonstrated reliable inference capabilities, but its performance lagged behind TensorRT in both precision configurations.
In numerical relativistic hydrodynamics simulations, the mapping of conserved variables to primitive variables often occurs tens of millions of times, underscoring the critical need for efficient inference strategies. While highly optimized multi-threaded CPU implementations, such as those achievable with the RePrimAnd library, can utilize up to 128 CPUs on platforms like Delta, GPU-based methods nonetheless offer a compelling alternative due to their massive parallelism. The inherent parallel architecture of GPUs enables the concurrent processing of large data batches, directly addressing the bottleneck posed by the numerous C2P calls required in large-scale simulations. Our results (Fig.4) demonstrate that TensorRT optimizations, particularly FP16 precision, further enhance GPU performance by enabling faster computations with minimal accuracy loss (Table 1). Furthermore, multi-GPU deployments, while not explored in this work, hold the potential to scale performance even further for the massive datasets characteristic of numerical relativity. While our single-GPU benchmarks do not demonstrate a decisive advantage over the single-core CPU implementation using RePrimAnd at the dataset scales tested, the observed trends in inference speed, coupled with ongoing advancements in GPU technology and software, suggest a promising outlook for GPU-based C2P solvers as dataset sizes and model complexities continue to increase.
IV Conclusions
@@
This work presents a novel method for accelerating C2P inversions in relativistic hydrodynamics simulations, focusing on hybrid piecewise polytropic and tabulated equations of state. We employ AI-driven techniques, specifically feedforward neural networks coupled with TensorRT optimization, to achieve significant performance improvements compared to traditional CPU solvers. Our results demonstrate that this approach offers a compelling alternative to computationally demanding iterative methods while maintaining a reasonably high accuracy.
Future research could explore several avenues for extending and refining this approach. This includes adapting the models to a broader class of equations of state and investigating alternative network architectures to further enhance performance. Continued exploitation of TensorRT, including additional optimization strategies and parallel implementations across multiple GPUs, offers the potential for even greater reductions in computational time for C2P inversions. These advancements could enable more efficient and detailed simulations of complex astrophysical systems. We anticipate that AI-driven methods, particularly those incorporating TensorRT, will be instrumental in advancing GRHD simulations, and the field of numerical relativity, in general. We release the software needed to validate and extend our findings at the following GitHub repository [33].
V Acknowledgements

### v2

@@ 2.3 Training approach
$\mathcal{L}(\theta)=\frac{1}{n}\sum_{i=1}^{n}(\hat{y}_{i}(\theta)-y_{i})^{2}+q\cdot\sum_{i=1}^{n}\mathrm{ReLU}(-\mathcal{N}^{-1}(\hat{y}_{i}(\theta)))\,,$ (5)
where $\hat{y}_{i}(\theta)$ represents the network’s estimation for feature $i$, $y_{i}$ is the corresponding target value, ReLU is the familiar rectified linear unit defined by $\mathrm{ReLU}(x)=\max(0,x)$, and $\mathcal{N}^{-1}(\cdot)$ represents an inverse normalization procedure based on the training data statistics. The penalty factor, $q$, was optimized for each model, with $q=150$ for NNC2PS and $q=350$ for NNC2PL. These values consistently suppressed negative pressure predictions on the test set. For the tabulated EOS model (NNC2P_Tabulated), the structure of the data precluded negative predictions, so a standard mean-squared error loss function was used.
@@
2.4 Inference Speed Tests
@@ 2.4.1 TorchScript Deployment
2.4.2 TensorRT Deployment
For TensorRT, we explored both FP32 (unquantized) and FP16-quantized engines, ultimately deciding not to pursue INT8 quantization due to accuracy degradation observed in initial tests. After extensive testing, we opted for dynamic engine building with a batch size determined by the total size of the expected dataset, as this approach provided the best balance between performance and flexibility for our hardware and model structure. It must be noted that constructing an optimal engine in TensorRT is a nuanced process, influenced by multiple factors including model architecture, hardware specifications, intended batch sizes during inference, and input data. Therefore, achieving the best results often involves iterative tuning and profiling to adapt the engine to the specific deployment environment and workload requirements. Below, we summarize the overall engine-building process we followed in detail:
@@ 2.4.2 TensorRT Deployment
  • Configuration and Optimization Profiles: The BuilderConfig was set with a 40 GB workspace memory limit, providing more than enough headroom for dynamic batch sizes while maintaining stable performance. We set up a dynamic optimization profile specifying minimum, optimal, and maximum batch sizes within 10 per cent margin of our typical usage, granting flexibility to handle both smaller and larger input volumes efficiently.
@@
For the actual inference speed test procedure, we implemented two distinct workflows on a single GPU for both approaches. The TorchScript-based approach allowed for a straightforward configuration, primarily requiring the definition of batch sizes and the pre-loading of data onto the GPU. It then used libtorch for efficient GPU deployment and batch execution.
In contrast, the TensorRT-based approach demanded several additional configurations. The model, after being converted into an optimized engine, was loaded using TensorRT’s C++ API. This included the manual pre-loading of input data into GPU memory before execution, and was followed by manual setup of input and output buffers for TensorRT’s executeV2 function, and careful management of CUDA resources. While this setup was more involved, it leveraged hardware-specific optimizations to deliver substantial gains in inference speed.
3 Results
3.1 Accuracy
@@
3.2 Inference Speed Analysis
@@
The inference performance of various methods was evaluated using a single NVIDIA A100 GPU for neural network models and a single-threaded CPU implementation of the traditional numerical method from the RePrimAnd library. The CPUs used in this study were dual AMD 64 core 2.45 GHz Milan processors on the Delta cluster, which can support up to 128 threads. Each configuration was tested across five dataset sizes, ranging from 25,000 to 1,000,000 data points, with ten inference runs conducted per configuration to ensure result stability and consistency.
The numerical method exhibited linear scaling of inference time with respect to the dataset size. In contrast, both TensorRT and TorchScript models generally maintained relatively stable inference times across the dataset sizes. Notably, the full-precision TensorRT engine for the smaller network, NNC2PS, showed a faster-than-expected processing time at certain intermediate dataset sizes, as observed in Fig. 4(a). This behavior may be attributed to favorable thread block utilization and the kernel selection mechanism of TensorRT for this particular network size. A more detailed profiling study is needed to fully elucidate the underlying cause. The accuracy characteristics of these models remained consistent, as indicated in Table 1.
Figure 4: Ideal scaling comparison of various C2P inversion methods under the assumption of perfect parallelization. (a) Projected inference time as a function of dataset size for a traditional numerical solver (RePrimAnd utilizing 128 CPU threads on a single node of the Delta cluster) and two neural network models (NNC2PS and NNC2PL) using TensorRT (FP32 and FP16 precision) and TorchScript across 8 NVIDIA A100 GPUs. (b) Projected inference speed comparison for a dataset of 8 million points, highlighting the significant scalability and efficiency gains achieved by TensorRT engines, particularly with FP16 optimization. The mixed-precision TensorRT engine for NNC2PS achieves approximately a 25-fold reduction in processing time compared to the numerical method, showcasing the potential for TensorRT-based methods to convincingly outperform traditional numerical solvers at scale.
The numerical method required significantly more time than the neural network-based approaches. On average, the numerical method took 103.8 ms to process 25,000 data points, with runtime scaling almost linearly to 3,490 ms for 1,000,000 data points. In contrast, the neural network models demonstrated substantially faster inference times. Specifically, the mixed-precision TensorRT engine built from NNC2PS required 7.92 ms for 25,000 data points and 8.54 ms for 1,000,000 data points. Its full-precision counterpart exhibited similar performance, with runtimes of 25.17 ms for 25,000 data points and 21.06 ms for 1,000,000 data points. The TorchScript variant showed slower performance, but still maintained sub-linear scaling, with runtimes averaging 72.79 ms for 25,000 points and 101.74 ms for 1,000,000 points.
A similar trend was observed for the NNC2PL models, with TensorRT engines consistently outperforming their TorchScript counterparts. The mixed-precision TensorRT engine for NNC2PL processed 25,000 data points in 8.32 ms and 1,000,000 points in 14.35 ms. In comparison, the full-precision TensorRT engine required 25.85 ms for 25,000 points and 23.87 ms for 1,000,000 points. The TorchScript model averaged 73.18 ms for 25,000 points and 102.04 ms for 1,000,000 points.
Figure 4 presents the ideal scaling achieved under the assumption of perfect parallelization, providing a theoretical performance benchmark. This scenario assumes optimal workload distribution, minimal communication overhead, and negligible synchronization delays, representing the upper bound of scalability. For the numerical method, the figure reflects the full computational capacity of a single CPU node on the Delta cluster, utilizing 128 threads. For the neural networks, it represents the use of 8 A100 GPUs within a single GPU node. Under these ideal conditions, the processing time of the numerical method per data point is projected to decrease by a factor of 128, allowing for the processing of 8 million points in approximately 218 ms (Fig. 4(b)). Similarly, all neural network methods are expected to achieve linear inference scaling with similar per-GPU efficiency. Under this scenario, TensorRT-based methods—particularly the mixed-precision engine for NNC2PS—show a 25-fold reduction in processing time for 8 million points compared to the numerical method running at full capacity on the CPU node. Furthermore, the scaling trend strongly favors TensorRT for even larger datasets.
The results presented above underscore the substantial performance gains achievable through the use of TensorRT-optimized neural networks, particularly in the context of conservative-to-primitive inversion in relativistic hydrodynamics simulations. By leveraging the parallel processing power of modern GPUs, these methods offer significant speedups compared to traditional CPU-based numerical approaches, even in large-scale simulations involving millions of data points. As demonstrated, TensorRT optimizations enable more efficient and scalable solutions, with the potential to dramatically reduce the computational cost of C2P operations. This work highlights the clear advantage of integrating ML-driven methods with GPU acceleration to address the computational challenges of high-throughput simulations. Moving forward, the next step is to incorporate these optimized approaches into full-scale hydrodynamics simulations, where their impact on both performance and scalability can be fully realized.
4 Conclusions
@@
This work introduces a novel ML-driven method for accelerating C2P inversions in relativistic hydrodynamics simulations, with a focus on hybrid piecewise polytropic and tabulated equations of state. By employing feedforward neural networks optimized with TensorRT, we achieve substantial performance improvements over traditional CPU solvers, offering a compelling alternative to computationally expensive iterative methods, while maintaining high accuracy. Our results demonstrate that the TensorRT-optimized neural networks can process large datasets significantly faster, achieving up to 25 times the inference speed of traditional methods.
Future work will explore several key directions to refine and expand this approach. First, adapting the models to handle a broader range of equations of state will improve the versatility of this method across different simulation contexts. Second, exploring alternative network architectures could further enhance both accuracy and inference speed. Additionally, continued optimization of TensorRT, including advanced parallelization strategies and scaling across multiple GPUs, promises even greater reductions in computational time, enabling simulations of larger and more complex astrophysical systems. These improvements will be critical for advancing high-resolution simulations in numerical relativistic hydrodynamics.
We believe that ML-driven methods, particularly those incorporating TensorRT optimization, will play an essential role in advancing the field of general relativistic hydrodynamics and numerical relativity more broadly. To facilitate further validation and extension of these findings, we have made the software developed for this study publicly available at the following GitHub repository [35].
5 Acknowledgements